OpenStudy (logo):

Can someone check if my answer is correct? What are the zeroes of y = x2 + 14x +40 ?

OpenStudy (logo):

y = 0 0 = x^2 + 14x +40 factors of 14 (a+b) = 14 = -10 -4 (a*b) = 40 = -10* -4 0 = x^2 - 10x - 4x + 40 x = -10 x = -4 y = -10^2 + 14*-10 + 40 y = 0 y = -4*-4 + 14*-4 + 40 y = 0

OpenStudy (steve816):

Yup! I got - 10 and -4 too Good job!

OpenStudy (logo):

@steve816 thx!

OpenStudy (steve816):

Welcome!